find JK if j(-3,-4) and k(-7,5)
Answer by jim_thompson5910(35256)

I'm assuming you want to find the length of segment JK. This is the same as finding the distance between the two points.


d = sqrt((x2-x1)^2+(y2-y1)^2)


d = sqrt((-7-(-3))^2+(5-(-4))^2)


d = sqrt((-7+3)^2+(5+4)^2)


d = sqrt((-4)^2+(9)^2)


d = sqrt(16+81)


d = sqrt(97)


So exact distance between the two points is sqrt(97) units.

This approximates to 9.8488578 units.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------